# Safe Reinforcement Learning on the Constraint Manifold: Theory and Applications

### v1

@@ VII-C Dynamic Mismatch in Air Hockey Simulation
The task requires controlling a KUKA iiwa14 robot to hit the randomly initialized puck into the opponent’s goal. The environment’s observation space contains the robot’s joint positions and velocities, the puck’s position and velocity, the relative position and velocity between the puck and the robot’s end-effector, and the action in the previous step. The total dimension of the observation space is 33. The action space of the RL agent is the desired joint velocity of the robot. A tracking controller is then applied to generate joint torques that track the desired joint position and velocity. The reward function is composed of three parts: one that rewards the end-effector for hitting the puck, one that expects the puck to move fast, and one that rewards for scoring. The safety specification is defined by 21 constraints, including 14 joint position constraints, five end-effector position constraints, and two elbow and wrist link constraints. Detailed descriptions of the environment and the hyperparameters are given in Appendix A-C.
Refer to captionRefer to caption
@@
To properly quantify the dynamics mismatch, we again use the velocity-controlled system as the nominal system for ATACOM and change the maximum velocity of the nominal system to be different percentages of the actual velocity limit. We trained the policy using SAC for 3 million steps and then ran the evaluation for 1000 episodes. Fig. 10 presents the results of 5 independent runs. We can observe that the performance (Discounted Return, Puck’s Velocity, and Success Rate) of ATACOM increases as the nominal system’s velocity increases. The performance gain is because the agent can hit the puck at a higher velocity, and the puck will score in fewer steps. However, this performance improvement comes at the cost of constraint violations. The ATACOM agent using an underestimated nominal dynamics model ($50\%,80\%,90\%$) generally has fewer violations than those using an overestimated model ($110\%,120\%,150\%$). The key reason is that the agent with an underestimated nominal dynamics model will always generate a feasible velocity command. In contrast, the agent with an overestimated nominal dynamics model will generate infeasible commands. The tracking controller will truncate the command to the velocity limit, which leads to unsafe behavior. Using a correct dynamics model, the agent achieves a good trade-off between performance and safety. In the experiment with an accurate nominal dynamics model, the violation rate is $0.1\%$: these violations are caused mainly by the low control frequency ($50$Hz) and the controller’s tracking error.
VII-D Real Robot Air Hockey
In the last experiment, we will show that ATACOM achieves safe exploration on the Real Robot Air Hockey task, enabling online RL in the real world. The real robot setup is illustrated in Figure 8d. The performance of the Robot Air Hockey Hitting task is susceptible to the robot’s and puck’s dynamics. The success rate of the trained agent is $87\%$ in simulation, yet the success rate drops to $12\%$ when the policy is deployed in the real world. The main reason behind the performance drop is that successful hitting requires the robot to strike the puck in a specific direction precisely, but the dynamic mismatch between the simulator and the real world leads to deviations in the tracking controllers and the puck movements. While the robot can still hit the puck at high speeds, the success rate is low. Therefore, training with real-world interactions is essential to obtain a high-performance agent. Unfortunately, training the agent from scratch on a real robot is expensive and time-consuming. One training experiment with 3 million steps will take more than 100 hours, including the time for resetting the task. Instead, we train the agent in a simulated air hockey environment until convergence and then continue online fine-tuning using real-world interactions.
@@ VII-D Real Robot Air Hockey
Figure. 11 shows the learning curve during the training process. The performance of the pre-trained agent is shown as a blue dashed line. The success rate of the pre-trained agent is $0.12$, and the average puck’s velocity is $0.97$m/s. We also compare the planning/optimization-based baseline as introduced in [13]. The baseline has a success rate of $0.3$ and an average hitting velocity of $1.23$m/s. Notably, the performance at 500 episodes is worse than the initial policy. This is because the value function trained in the simulation does not reflect the performance of the policy in the real world, and the value function needs more iterations to reconstruct the landscape based on real-world dynamics. Since the agent has already acquired some high-performance data in the initial data collection phase, the value function will converge faster than if it is trained from scratch. The success rate improved significantly to $0.71$ at the evaluation of 3000 episodes, and the puck’s speed remains comparable at $0.97$m/s. The maximum violation throughout the evaluation is close to zero, as shown in the lower-right plot of Figure 11. Although the planning baseline is able to hit the puck at a higher speed, building a reactive and adaptive planner for high-speed dynamic motion is still challenging. The goal of the learning agent is focused primarily on scoring, which sacrifices hitting speed for better accuracy. We can also observe it from the last two steps in the learning curve. The success rate keeps increasing while the hitting velocity decreases slightly.
We show the success rate when the puck is initialized at different table regions in Figure 12. The planning/optimization-based solution is shown in the leftmost plot. The three plots on the right show the success rates at Episodes $0,1500$, and $3000$. Due to the uneven air flow, the puck drifts heavier on the right side of the table than on the left. Therefore, the baseline agent performs worse on the right side as the planner does not adapt to the changes in the puck’s motion. After the training, we can observe that the RL agent successfully adapted its behavior, and the success rate is significantly improved. However, the puck’s speed is still lower than the baseline as the baseline is optimized for the maximum speed, while the RL agent only contains a small bonus for the high-speed puck. The RL agent needs to make a trade-off between high-speed motion and accuracy. Although other specialized solutions to solve the Air Hockey hitting task exist [76], to the best of our knowledge, our method is the first approach able to learn and improve the performance of the real robot while ensuring safety at every timestep. We argue that the RL agent could, in principle, outperform these baselines using more effective learning strategies and fine-tuning the reward design.
@@ VIII-A Connections to Control Barrier Functions
A safe controller is often obtained from a quadratic programming problem that finds the control input closest to the nominal one while satisfying the CBF constraint (18).
We found a close connection between ATACOM and CBF in the following sense: consider the 1-dimensional safety constraint and the following equality constraint
$k(\bm{s})+\mu=0$
with $\mu\geq 0$. We rearrange the equality constraint and get
$\mu(\bm{s})=-k(\bm{s}),\quad\text{and}\quad\dot{\mu}(\bm{s})=-\dot{k}(\bm{s})$
From the slack variable dynamics (8), we have
$\dot{\mu}(\bm{s})=\alpha(\mu(\bm{s}))u_{\mu}\geq u_{\mu}^{-}\alpha(\mu(\bm{s}))=-\alpha^{\prime}(\mu(\bm{s}))$
with $\alpha^{\prime}(\cdot)=-u_{\mu}^{-}\alpha(\cdot)$. The inequality holds since $\alpha(\cdot)\geq 0$. Since $u_{\mu}^{-}<0$, $\alpha^{\prime}$ is also a function of class $\mathcal{K}$. Compared to the CBF constraint (18), introducing a slack dynamics function is equivalent to constructing a CBF constraint.
The major difference between ATACOM and CBF is that ATACOM obtains the safe action by constructing a set of basis vectors while CBF-base approach requires an additional optimization step to find the feasible action. Solving this optimization step numerically leads to additional computation costs. Instead, ATACOM only requires matrix decomposition and multiplication at each time step. We can construct the slack dynamics function from the CBF and vice versa by $h(\cdot)=-k(\cdot)$ and $\alpha^{\prime}(\cdot)=-u_{\mu}^{-}\alpha(\cdot)$. Another difference between the two methods is when no feasible action exists to satisfy the constraint. CBF-based solution does not provide a alternative solution while ATACOM provides a least square solution.
@@ VIII-B Limitations
  • In our analysis, we assume a feasible action exists such that Equation (11) is solvable. The limit of the action is not considered. It is generally a challenging problem to analyze the stability with control input limits. Constructing a safety constraint considering the actuation limit is also a challenging problem. Learning-based approaches to construct a CBF with control input limit is an active topic [77, 78].
  • Our approach only considers single-step constraints. It does not consider long-term safety. For example, the robot may get stuck in the middle of multiple objects in the 2D Moving Obstacle environment. This type of safety specification can be addressed by restricting a cumulative cost in SafeRL algorithms.
@@ VIII-C Conclusions
In this paper, we provided a theoretical foundation, a set of extensions, and through hyperparameters studies for ATACOM algorithm. We showed that safety constraints can be constructed as a manifold. By exploiting the geometry of the tangent space, we can generate a safe action space, allowing learning agents to sample arbitrary actions while ensuring safety. The theoretical analysis demonstrated the existence of a region of attraction around the constraint manifold, guaranteeing the system converges to the manifold. We demonstrated in a real-world experiment of the robot air hockey task that our approach can learn a safe policy in a high-dimensional task with complex safety constraints.
@@ IX Acknowledgments
Appendix A Hyperparameters for the Experiments
A-A Hyperparameters for the 2D Static Environment
@@ A-B Hyperparameters for the 2D Dynamic Environment
A-C Experiment setup for the Robot Air Hockey Task
The observation space for the robot air hockey task is
${\bm{u}}=r_{u}{\bm{u}}_{\text{sample}}+(1-r_{u}){\bm{u}}_{\text{prev}}$
where $r_{u}$ is the smoothing ratio. The robot is controlled by torque in $1000$Hz and the RL action is sampled in $50$Hz. We applied a trajectory interpolator to convert the desired joint velocity to a chunk of trajectories in $1000$Hz. Then a PD controller is used to track the trajectory. The constraints are defined as
${\bm{q}}_{i,l}<{\bm{q}}_{i}<{\bm{q}}_{i,u},\quad i=1,\cdots,7$
$x_{l}<x_{\text{ee}},\quad y_{l}<y_{\text{ee}}<y_{u},\quad z_{l}<z_{\text{ee}}<z_{u}$
$z_{l}^{\prime}<z_{\text{wrist}}\quad z_{l}^{\prime}<z_{\text{elbow}}$
where ${\bm{q}}_{i}$, ${\bm{q}}_{i,l}$ and ${\bm{q}}_{i,u}$ are the position, lower and upper limits of the joint $i$. $[x_{\text{ee}},y_{\text{ee}},z_{\text{ee}}]^{\intercal}=\text{FK}({\bm{q}})$ is the end-effector position obtained from the forward kinematics. $z_{\text{wrist}}$ and $z_{\text{elbow}}$ are the $z$-axis position of the wrist and elbow. The total number of constraints is 21.
The reward function is defined as follows
$r(\bm{s},{\bm{u}}_{s})$ $=r_{\text{pe}}+r_{\text{pv}}+r_{\text{pg}},$
$r_{\text{pv}}$ $=1.5\dot{x}_{\text{puck}},\text{ if }x_{\text{puck}}>0$
where $r_{\text{pe}}$ defines the reward that encourages the end-effector to approach the puck, $d_{\text{pe}}$ is the minimum distance between the puck and the end-effector in the episode. $r_{\text{pv}}$ is the reward that encourages the puck to move at a higher speed. $r_{\text{pg}}$ is the reward at the final step of the episode encouraging the puck to reach the goal.
The parameters for the RL training are shown in Table V.
@@ A-C Experiment setup for the Robot Air Hockey Task
A-D Experiment Setup for the Real Robot Air Hockey Task
In this experiment, the pre-training in simulation is conducted with the same hyperparameters as in Table V. Here we list the hyperparameters for the real robot experiment.

### v2

@@ VII-D Dynamic Mismatch in Air Hockey Simulation
The task requires controlling a KUKA iiwa14 robot to hit the randomly initialized puck into the opponent’s goal. The observation includes the robot’s joint positions (7-dim) and velocities (7-dim), the puck’s position (3-dim) and velocity (3-dim), the relative position (3-dim) and velocity (3-dim) between the puck and the robot’s end-effector, and the action (7-dim) in the previous step. The total dimension of the observation space is 33. The action space of the RL agent is the desired joint velocity (7-dim) of the robot. A feedforward controller is then applied to generate joint torques tracking the desired joint position and velocity. The reward function is composed of three parts: one that rewards the end-effector for hitting the puck, one that expects the puck to move fast, and one that rewards for scoring. Safety is ensured by imposing constraints in both the joint space and Cartesian space. In the joint space, 14 constraints limit the joint positions within their allowable ranges. For the Cartesian space, two constraints define the upper and lower bounds on the end-effector’s height. Additionally, three constraints (two in the $Y$-direction and one in the $X$-direction) prevent the end-effector from colliding with the table’s boundaries. To further ensure safety, two additional constraints restrict the height of the robot’s elbow and shoulder, avoiding potential collisions with the table. A detailed description of the environment and hyperparameters can be found in Appendix A.
@@
To properly quantify the dynamics mismatch, we again use the velocity-controlled system as the nominal system for ATACOM and change the maximum velocity of the nominal system to be different percentages of the actual velocity limit. We trained the policy using SAC for 3 million steps and then ran the evaluation for 1000 episodes. Fig. 10 presents the results of 5 independent runs. We can observe that the performance (Discounted Return, Puck’s Velocity, and Success Rate) of ATACOM increases as the nominal system’s velocity increases. The performance gain is because the agent can hit the puck at a higher velocity, and the puck will score in fewer steps. However, this performance improvement comes at the cost of constraint violations. The ATACOM agent using an underestimated nominal dynamics model ($50\%,80\%,90\%$) generally has fewer violations than those using an overestimated model ($110\%,120\%,150\%$). The key reason is that the agent with an underestimated nominal dynamics model will always generate a feasible velocity command. In contrast, the agent with an overestimated nominal dynamics model will generate infeasible commands. The tracking controller will truncate the command to the velocity limit, which leads to unsafe behavior. Using a correct dynamics model, the agent achieves a good trade-off between performance and safety. In the experiment with an accurate nominal model, the violation rate is $0.1\%$: these violations are caused mainly by the low control frequency ($50$Hz) and the controller’s tracking error.
VII-E Real Robot Air Hockey
In the last experiment, we will show that ATACOM achieves safe exploration on the Real Robot Air Hockey task, enabling online RL in the real world. The real robot setup is illustrated in Fig. 7e. The performance of the Robot Air Hockey Hitting task is susceptible to the robot’s and puck’s dynamics. The success rate of the trained agent is $87\%$ in simulation, yet the success rate drops to $12\%$ when the policy is deployed in the real world. The main reason behind the performance drop is that successful hitting requires the robot to strike the puck in a specific direction precisely, but the dynamic mismatch between the simulator and the real world leads to deviations in the tracking controllers and the puck movements. While the robot can still hit the puck at high speeds, the success rate is low. Therefore, training with real-world interactions is essential to obtain a high-performance agent. Unfortunately, training the agent from scratch on a real robot is expensive and time-consuming. One training experiment with 3 million steps will take more than 100 hours, including the time for resetting the task. Instead, we train the agent in a simulated air hockey environment until convergence and then continue online fine-tuning using real-world interactions.
@@ VII-E Real Robot Air Hockey
Fig. 11 shows the learning curve during the training process. The performance of the pre-trained agent is shown as a blue dashed line. The success rate of the pre-trained agent is $0.12$, and the average puck’s velocity is $0.97$m/s. We also compare the planning/optimization-based baseline as introduced in [82]. The baseline has a success rate of $0.3$ and an average hitting velocity of $1.23$m/s. Notably, the performance at 500 episodes is worse than the initial policy. This is because the value function trained in the simulation does not reflect the performance of the policy in the real world, and the value function needs more iterations to reconstruct the landscape based on real-world dynamics. Since the agent has already acquired some high-performance data in the initial data collection phase, the value function will converge faster than if it is trained from scratch. The success rate improved significantly to $0.71$ at the evaluation of 3000 episodes, and the puck’s speed remains comparable at $0.97$m/s. The maximum violation throughout the evaluation is close to zero, as shown in the lower-right plot of Fig. 11. Although the planning baseline is able to hit the puck at a higher speed, building a reactive and adaptive planner for high-speed dynamic motion is still challenging. The goal of the learning agent is focused primarily on scoring, which sacrifices hitting speed for better accuracy. We can also observe it from the last two steps in the learning curve. The success rate keeps increasing while the hitting velocity decreases slightly.
We show the success rate when the puck is initialized at different table regions in Fig. 12. The planning/optimization-based solution is shown in the leftmost plot. The three plots on the right show the success rates at Episodes $0,1500$, and $3000$. Due to the uneven air flow, the puck drifts heavier on the right side of the table than on the left. Therefore, the baseline agent performs worse on the right side as the planner does not adapt to the changes in the puck’s motion. After the training, we can observe that the RL agent successfully adapted its behavior, and the success rate is significantly improved. However, the puck’s speed is still lower than the baseline as the baseline is optimized for the maximum speed, while the RL agent only contains a small bonus for the high-speed puck. The RL agent needs to make a trade-off between high-speed motion and accuracy. Although other specialized solutions to solve the Air Hockey hitting task exist [83], to the best of our knowledge, our method is the first approach able to learn and improve the performance of the real robot while ensuring safety at every timestep. We argue that the RL agent could, in principle, outperform these baselines using more effective learning strategies and fine-tuning the reward design.
@@ VIII-A Connections to Control Barrier Functions
A safe controller is often obtained from a quadratic programming problem that finds the control input closest to the nominal one while satisfying the CBF constraint (19). Let $h(\bm{s})=-k(\bm{s})=\bm{\mu},\alpha(h(\bm{s}))=-{\bm{u}}_{\mu}\alpha^{\prime}(h(\bm{s}))$, the CBF constraint results in a formulation similar to ATACOM
$\underbrace{{\bm{J}}_{k}\bm{f}}_{{\bm{\psi}}}+\underbrace{[{\bm{J}}_{k}\bm{G}\quad\text{diag}(\alpha^{\prime}(\bm{\mu})]}_{{\bm{J}}_{u}}\begin{bmatrix}{\bm{u}}_{s}\\ {\bm{u}}_{\mu}\end{bmatrix}\leq{\bm{0}}\quad\exists{\bm{u}}_{s}\in\mathcal{U}$
while ATACOM considers the equality constraint (11).
However, the methods differ in how these constraints are solved. ATACOM introduces a drift compensation term $-{\bm{J}}_{u}^{\dagger}{\bm{\psi}}$ that solves the least-norm problem
$\min\|[{\bm{u}}_{s}\quad{\bm{u}}_{\mu}]^{\intercal}\|^{2}\quad\mathrm{s.t.}\;{\bm{\psi}}+{\bm{J}}_{u}\begin{bmatrix}{\bm{u}}_{s}\\ {\bm{u}}_{\mu}\end{bmatrix}={\bm{0}},{\bm{u}}_{s}\in\mathcal{U}$ (20)
The external command for ATACOM is applied in the tangent space, requiring only matrix decomposition and multiplication. Instead, a typical CBF solution involves formulating a Least-square problem w.r.t the reference action ${\bm{u}}_{r}$
$\min\|{\bm{u}}_{s}-{\bm{u}}_{r}||^{2}\quad\mathrm{s.t.}\quad{\bm{\psi}}+{\bm{J}}_{k}{\bm{u}}_{s}\leq{\bm{0}}$
Here, the external command is modified through an implicit optimization process, often involving iterative methods, such as the Active Set Method or the Interior Point Method [84].
A key advantage of ATACOM is that it constructs a direct mapping (12), allowing the gradient $\nabla_{{\bm{u}}}{\bm{u}}_{s}$ to be computed straightforwardly. In contrast, CBF must rely on differentiable solvers, which can be computationally expensive. Gradient propagation is crucial when jointly learning constraints and policy [85] or for constraint-agnostic task learning [48] Additionally, when no feasible action exists to satisfy the constraint. CBF-based solution fails to find an alternative solution while ATACOM provides a least square solution to (20).
@@ VIII-C Conclusions
In this paper, we provided a theoretical foundation, a set of extensions, and through hyperparameters studies for ATACOM algorithm. We showed that safety constraints can be constructed as a manifold. By exploiting the geometry of the tangent space, we can generate a safe action space, allowing learning agents to sample arbitrary actions while ensuring safety. The theoretical analysis demonstrated the existence of a region of attraction around the constraint manifold, guaranteeing the system converges to the manifold. Furthermore, we also analyze the safety of the system under disturbances will ensure bounded constraint violations. We demonstrated in a real-world experiment of the robot air hockey task that our approach can learn a safe policy in a high-dimensional task with complex safety constraints.
@@ IX Acknowledgments
Appendix A Experiment setup for the Robot Air Hockey Task
The observation space for the robot air hockey task is
${\bm{u}}=r_{u}{\bm{u}}_{\text{sample}}+(1-r_{u}){\bm{u}}_{\text{prev}}$
where $r_{u}$ is the smoothing ratio. The robot is controlled by torque in $1000$Hz and the RL action is sampled in $50$Hz. We applied a trajectory interpolator to convert the desired joint velocity to a chunk of trajectories in $1000$Hz. Then a PD controller is used to track the trajectory. The constraints are
$q_{i,l}<q_{i}<q_{i,u},\quad i=1,\cdots,7$
$x_{l}<x_{\text{ee}},\quad y_{l}<y_{\text{ee}}<y_{u},\quad z_{l}<z_{\text{ee}}<z_{u}$
$z_{l}^{\prime}<z_{\text{wrist}}\quad z_{l}^{\prime}<z_{\text{elbow}}$
where $q_{i}$, $q_{i,l}$ and $q_{i,u}$ are the position, lower and upper limits of the joint $i$. $[x_{\text{ee}},y_{\text{ee}},z_{\text{ee}}]^{\intercal}=\text{FK}({\bm{q}})$ is the end-effector position obtained from the forward kinematics. $z_{\text{wrist}}$ and $z_{\text{elbow}}$ are the $z$-axis position of the wrist and elbow. The total number of constraints is 21.
The reward function is defined as follows
$r(\bm{s},{\bm{u}}_{s})$ $=r_{\text{pe}}+r_{\text{pv}}+r_{\text{pg}},$
$r_{\text{pv}}$ $=1.5\dot{x}_{\text{puck}},\text{ if }x_{\text{puck}}>0$
where $r_{\text{pe}}$ defines the reward that encourages the end-effector to approach the puck, $d_{\text{pe}}$ is the minimum distance between the puck and the end-effector in the episode. $r_{\text{pv}}$ is the reward that encourages the puck to move at a higher speed. $r_{\text{pg}}$ is the reward at the final step of the episode, encouraging the puck to reach the goal. The parameters for the RL training are shown in Table V.
Appendix B Hyperparameters
@@
In the real-world experiment, the pre-training in simulation is conducted with the same hyperparameters as in Table V. Here we list the hyperparameters for the real robot experiment.